OpenStudy (anonymous):

Please Help Will Fan And Medal! :) What is the perimeter of a rectangle with a length of 7 3/4 inches and a width of 4 1/3

OpenStudy (anonymous):

What do u think

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

Sorry I can't see your options u have to solve it out

OpenStudy (anonymous):

Oh Sorry I Think its 11 4/7

OpenStudy (anonymous):

In your option do u have 241/10

OpenStudy (anonymous):

Meaning 24 1/10

OpenStudy (anonymous):

Nope These Are My Options A. 12 1/12 B. 24 1/3 C. 11 4/7 D 24 1/6

OpenStudy (anonymous):

Well the formula for perimeter of a rectangle is P=2(l+w). First convert the mixed fractions into improper fractions.

OpenStudy (anonymous):

Convert this to a mixed fraction 145/6

OpenStudy (anonymous):

89/6??

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Ok hold on do u have calculator with u

OpenStudy (anonymous):

Yeah Why?

OpenStudy (anonymous):

Ok first multiply 7 and 4

OpenStudy (anonymous):

from 7 3/4

OpenStudy (anonymous):

Okay if u are using Casio Type out 145/6 Shift And push the button that identifies S-D

OpenStudy (anonymous):

28

OpenStudy (anonymous):

(first multiply the denominator and whole number)

OpenStudy (anonymous):

Correct 28, now add that to the numerator

OpenStudy (anonymous):

(3)

OpenStudy (anonymous):

31

OpenStudy (anonymous):

So 7 3/4= 31/4 Do the same to 4 1/3

OpenStudy (anonymous):

13

OpenStudy (anonymous):

@Lovelt help him in covering 145/6 To a mixed fraction

OpenStudy (anonymous):

Im A Girl -.-

OpenStudy (anonymous):

Sorry her

OpenStudy (anonymous):

Am so so sorry about mistaking your gender

OpenStudy (anonymous):